# EZ-CLIP: Efficient zero-shot video action recognition

### 13 Dec 2023 v1

@@ 1 Introduction
Figure 1: Effectiveness of EZ-CLIP: (Left) Performance comparison on UCF-101 dataset for zero-shot evaluation. The proposed method outperforms existing works while requiring fewer tunable parameters and GFLOPS with higher throughput (TP). Bubble size indicates GFLOPS during inference. (Right) Performance comparison on Something-something-v2 (SSv2) for base-to-novel evaluation. EZ-CLIP achieves significant improvement with fewer tunable parameters over existing works on this challenging dataset where motion plays a critical role.
@@ 4.1 Evaluation and comparisons
Base-to-Novel: To assess the generalization capabilities of EZ-CLIP towards novel classes, we evaluate it in a base-to-novel setting Rasheed et al. (2023). We begin with a dataset $D_{S}$ with labels $Y_{S}=\{y_{i}\}_{i=0}^{k}$, which is partition into two categories: the base classes $Y_{B}$ and the novel classes $Y_{N}$. This partition ensures that $Y_{B}\cup Y_{N}=Y_{S}$ and $Y_{B}\cap Y_{N}=\emptyset$. The model is trained on the base classes and then subjected to evaluation on both the base and novel classes. The evaluation is shown in Table 2 for four diverse datasets: K-400, HMDB-51, UCF-101, and SSv2. EZ-CLIP consistently outperforms existing models, showcasing significant improvements across all datasets. The most significant improvement is observed in SSv2 dataset, where it achieves a substantial +37.8% gain in base class accuracy and an impressive +8.6% improvement in the Novel class of SSv2.
@@
Table 4: Ablation study: We evaluate the effectiveness of temporal visual prompting (TVP) and motion loss using base-to-novel setup on Kinetics-400, UCF-101, HMDB-51 and SSv2.
@@ 4.2 Ablations
Impact of temporal visual prompting: We include temporal visual prompting in the base model to study its impact. We observe in Tables 4, 10 and 11 that temporal visual prompting consistently improves accuracy over the base model for both base and novel classes and across all the datasets. In addition, we also observe that the improvement is more significant on SSv2 dataset in comparison with HMDB-51 and UCF-101 datasets, which signifies its temporal modeling capability.
Impact of Motion Loss: Adding the proposed motion loss as a constraint during base model training consistently enhances performance, albeit to a lesser extent than observed with temporal prompting (Tables 4, 10, 11). This improvement is attributed to the model’s limited ability to learn temporal aspects, given the absence of temporal prompts. More detailed insights into the impact of Motion Loss on video frame embeddings are provided in Section (A.9).
@@ 4.3 Discussion and analysis
Template prompt vs action class description: Tables 5,9 and 12 provides a detailed comparison of the model’s performance with and without LLM-generated descriptions on two datasets: UCF-101 and HMDB-51. Our findings reveal that incorporating LLM-generated action class descriptions leads to improved text embedding generalization and enhanced model accuracy.
Table 5: Descriptive prompts: Impact of LLM-generated action class descriptions as prompts.

### 19 Jan 2024 v2

@@ 1 Introduction
Figure 1: Effectiveness of EZ-CLIP: Performance comparison on UCF-101 dataset for zero-shot evaluation. The proposed method outperforms existing works while requiring fewer tunable parameters and GFLOPS with higher throughput (TP). Bubble size indicates GFLOPS during inference. (Right) Performance comparison on Something-something-v2 (SSv2) for base-to-novel evaluation. EZ-CLIP achieves significant improvement with fewer tunable parameters over existing works on this challenging dataset where motion plays a critical role.
@@ 4.1 Evaluation and comparisons
EZ-CLIP(ViT-16) 73.1 60.6 66.3 77.0 58.2 66.3 94.4 77.9 85.4 16.6 13.3 14.8
Table 3: Base to novel generalization all sample: Evaluation results for Base-to-Novel setting using all samples in the base class.
Kinetics-400 HMDB-51 UCF-101 SSv2
Method Base Novel HM Base Novel HM Base Novel HM Base Novel HM
@@
Base-to-Novel: To assess the generalization capabilities of EZ-CLIP towards novel classes, we evaluate it in a base-to-novel setting Rasheed et al. (2023). We begin with a dataset $D_{S}$ with labels $Y_{S}=\{y_{i}\}_{i=0}^{k}$, which is partitioned into two categories: the base classes $Y_{B}$ and the novel classes $Y_{N}$. This partition ensures that $Y_{B}\cup Y_{N}=Y_{S}$ and $Y_{B}\cap Y_{N}=\emptyset$. The model is trained on the base classes using 16 samples per base class. The evaluation is shown in Table 2 for four diverse datasets: K-400, HMDB-51, UCF-101, and SSv2. EZ-CLIP consistently outperforms existing models, showcasing significant improvements datasets HMDB-51, UCF-101, and SSv2.
Base-to-Novel trained on all samples: we also trained the model for all samples in the base class to showcase the full capacity of EZ-CLIP. The evaluation is shown in Table 3 for four diverse datasets: K-400, HMDB-51, UCF-101, and SSv2. EZ-CLIP consistently outperforms existing models, showcasing significant improvements across all datasets.
@@
Table 5: Ablation study: We evaluate the effectiveness of temporal visual prompting (TVP) and motion loss using the base-to-novel setup on Kinetics-400, UCF-101, HMDB-51, and SSv2, using all samples from the base classes.
@@ 4.2 Ablations
Impact of temporal visual prompting: We include temporal visual prompting in the base model to study its impact. We observe in Tables 5, 11 and 12 that temporal visual prompting consistently improves accuracy over the base model for both base and novel classes and across all the datasets. In addition, we also observe that the improvement is more significant on SSv2 dataset in comparison with HMDB-51 and UCF-101 datasets, which signifies its temporal modeling capability.
Impact of Motion Loss: Adding the proposed motion loss as a constraint during base model training consistently enhances performance, albeit to a lesser extent than observed with temporal prompting (Tables 5, 11, 12). This improvement is attributed to the model’s limited ability to learn temporal aspects, given the absence of temporal prompts. More detailed insights into the impact of Motion Loss on video frame embeddings are provided in Section (A.9).
@@ 4.3 Discussion and analysis
Template prompt vs action class description: Tables 6,10 and 13 provides a detailed comparison of the model’s performance with and without LLM-generated descriptions on two datasets: UCF-101 and HMDB-51. Our findings reveal that incorporating LLM-generated action class descriptions leads to improved text embedding generalization and enhanced model accuracy.
Table 6: Descriptive prompts: Impact of LLM-generated action class descriptions as prompts, utilizing all samples from the base classes.